OpenStudy (anonymous):

Can Someone Also Help Me Understand How I Got The Wrong Answer To This Question I Got b=5 I Dont Know How BUt The Question Is b-9+6b=30

OpenStudy (anonymous):

I'm not sure how you got b = 5, if you show the steps you took to finding that solution I could find the mistake you made. But otherwise, the solution is as follows: first combine like terms b - 9 + 6b = 7b - 9 because we can combine the two terms that have a b in them so 7b - 9 =30 then solve for b 7b = 39 b = 39/7

OpenStudy (anonymous):

I Got 39/7 = 5.571428571428571

OpenStudy (anonymous):

typically you can leave your solution in the form of a fraction, because nobody cares about a number with 20 decimal places in it, so 39/7 will suffice

OpenStudy (anonymous):

Ok Thanks I Knew My Mom was wrong!

OpenStudy (anonymous):

haha be careful how often you think that :)